# Holistic Evaluation for Interleaved Text-and-Image Generation

### v2

@@ Abstract
Minqian Liu  Zhiyang Xu  Zihao Lin  Trevor Ashby  Joy Rimchala Jiaxin ZhangLifu Huang Virginia Tech  Intuit AI Research {minqianliu, zhiyangx, zihaol, trevorashby, lifuh}@vt.edu {joy_rimchala, jiaxin_zhang}@intuit.com
@@ 2 Related Work
Evaluating multimodal generation tasks presents unique challenges due to the inherent complexity of assessing both textual and visual components simultaneously. Existing metrics for text generation, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and LLM-based evaluators Zhong et al. (2022); Liu et al. (2023b, 2024), fall short when applied to multimodal outputs as they fail to capture the visual quality and coherence with textual content. Similarly, visual generation metrics like FID Heusel et al. (2017) and IS Salimans et al. (2016) are inadequate for evaluating the textual elements accompanying the images. To address this, recent studies have employed multimodal metrics Zhang et al. (2023b); Ku et al. (2023), such as CLIPScore Hessel et al. (2021), which leverages the alignment capabilities of the CLIP model to measure the similarity between generated images and their corresponding textual descriptions. However, CLIPScore can only measure the alignment between text and images, which is not sufficient to evaluate the quality of generated output comprehensively. Moreover, human evaluations, although more reliable, are resource-intensive and cannot be scalable. In terms of evaluation benchmarks in multimodal learning, existing works mostly focus on evaluating the tasks with single-modality output Fu et al. (2024); Li et al. (2024); Lu et al. (2024); Wang et al. (2024), such as conditional text-to-image generation Chen et al. (2024); Ku et al. (2024), where the primary focus is solely the quality of generated images. Our InterleavedBench benchmark introduces a novel approach to evaluate interleaved text-and-image generation by incorporating multiple aspects of quality assessment, thus providing a more nuanced and holistic evaluation framework.
@@ 5 Experiments
Table 2: Automatic evaluation results of existing interleaved generation models on InterleavedBench using InterleavedEval. TIC is the abbreviation for ’Text-Image Coherence’. The best results are highlighted in bold.
@@ 5.2 Main Results
We show the main results of using InterleavedEval to conduct the fine-grained evaluation for various baseline approaches on InterleavedBench in Table 2. The baselines in the upper part are the integrated and open-sourced models while the baselines in the lower part are the pipeline models where the LMMs are proprietary. From Table 2, we observe that: First, the pipeline models consistently outperform the integrated models on all evaluation aspects by a significant margin, where GPT-4o + DALL·E 3 achieves the best performance on helpfulness and the average score of all the aspects. Second, the pipeline models achieve significantly good performance on text quality since Gemini and GPT-4o have strong text generation capabilities. Also, the generated visual prompts are generally coherent with the text content and they are directly fed into the image generation model, so the performance on text-image coherence of pipeline models is also remarkable. Third, we observe that the common errors of integrated models include the output text and/or images being empty, in poor quality, or having severe duplication. This is probably due to their weak instruction-following abilities. Fourth, image coherence is the most challenging aspect for the pipeline models. This is because the image generation model cannot take the images in the input context or previously generated images as conditions. Thus, the generated images do not have strong coherence. We include more qualitative analysis to illustrate these observations in Section 6.
@@ Results
To validate the effectiveness of our proposed metric, we conduct a correlation analysis by comparing the evaluation results from automatic metrics with our human evaluation results. Since the baseline metrics only predict an overall score for each instance, we use the same set of evaluation scores to compare against the human rating on each aspect separately. For InterleavedEval, we compare evaluation scores with the human rating on corresponding aspects. Since most baselines require a gold reference, we use the context-based subset, where each instance has an associated reference output, to compute the correlation. From Table 4, our InterleavedEval consistently outperforms previous metrics by a significant margin in every aspect. Our metric has a particularly higher correlation on text quality, which is because text quality is relatively easier to evaluate with large language models like GPT-4o Zheng et al. (2023b). Our metric achieves the lowest correlation on perceptual quality. The plausible reason is that GPT-4’s perceptual recognition capability is still not strong enough to accurately detect visual artifacts or unnatural disruptions in the images Fu et al. (2024). We also find that baseline metrics generally achieve poorer correlation, e.g., most metrics except for BERTScore almost do not have any correlation with helpfulness. BERTScore achieves the best correlation on helpfulness among baseline metrics, which indicates that text quality could be a good indicator of whether the overall interleaved content is helpful.
@@ 6 Discussions
Figure 6: Radar figure of evaluation results on each evaluation aspect for each task.
@@
We conduct a qualitative analysis of benchmarked models in Figure 5 and have the following observations: (1) while GILL can generate images with reasonable quality, the generated text and images are typically not coherent with the instruction and context. In the example in the first row, the generated text is totally irrelevant to the task, while the image is also inconsistent with input images. (2) EMU-2 can often generate text that is relevant to the task, but the quality is not good enough. In the example in the second row, it repeatedly says “soak the fabric in water” but does not contain other useful content. Another weakness of EMU-2 is its poor conditional image generation capability, where generated images have obvious visual distortions and could be duplicated with input images. (3) On the other hand, the pipeline models can generally better follow the instructions and generate text and images in higher quality. Nevertheless, they still occasionally have some drawbacks. For Gemini+SDXL, some of the generated images (e.g., the first output image in the second example) still have obvious defects. For GPT-4+DALLE3, the style of generated images can be dramatically different from input images, as DALLE3 is prone to generate images in cartoon or dramatic styles. (4) Maintaining image coherence, i.e., the coherence of style and entities across images, is still very challenging for most models. In the third example, for the pipeline models, the same character has a very different appearance across the images, which makes the content inconsistent. (5) For the instances on the context-free subset, the integrated baselines have significantly worse performance, where they only generate one image with extremely poor quality. We hypothesize the reason to be those models cannot truly understand and follow the instructions. To sum up, our qualitative analysis indicates there is still significant room for improvement in interleaved generation.
@@ Appendix B More Details on Evaluation
B.1 Human Evaluation
More Details on Human Evaluation Setup
Inter-Annotator Agreement
We show the IAA of our human evaluation in Table 6. While our human evaluation did not achieve significantly high agreement, we argue that the evaluation of interleaved generation is still quite subjective, open-ended, and challenging, even with our carefully designed human evaluation aspects and guidelines.
Appendix C Additional Experiment Results

### v3

@@ Abstract
Minqian Liu  Zhiyang Xu  Zihao Lin  Trevor Ashby Joy RimchalaJiaxin ZhangLifu Huang♠,♣ Virginia Tech  Intuit AI Research  University of California, Davis {minqianliu, zhiyangx, zihaol, trevorashby, lifuh}@vt.edu {joy_rimchala, jiaxin_zhang}@intuit.com
@@ 2 Related Work
Evaluating multimodal generation tasks presents unique challenges due to the inherent complexity of assessing both textual and visual components simultaneously. Existing metrics for text generation, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and LLM-based evaluators Zhong et al. (2022); Liu et al. (2023b, 2024), fall short when applied to multimodal outputs as they fail to capture the visual quality and coherence with textual content. Similarly, visual generation metrics like FID Heusel et al. (2017) and IS Salimans et al. (2016) are inadequate for evaluating the textual elements accompanying the images. To address this, recent studies have employed multimodal metrics Zhang et al. (2023b); Ku et al. (2023), such as CLIPScore Hessel et al. (2021), which leverages the alignment capabilities of the CLIP model to measure the similarity between generated images and their corresponding textual descriptions. However, CLIPScore can only measure the alignment between text and images, which is not sufficient to evaluate the quality of generated output comprehensively. Moreover, human evaluations, although more reliable, are resource-intensive and cannot be scalable. In terms of evaluation benchmarks in multimodal learning, existing works mostly focus on evaluating the tasks with single-modality output Fu et al. (2024b); Li et al. (2024a); Lu et al. (2024); Wang et al. (2024); Fu et al. (2024a), such as conditional text-to-image generation Chen et al. (2024); Ku et al. (2024), where the primary focus is solely the quality of generated images. Our InterleavedBench benchmark introduces a novel approach to evaluate interleaved text-and-image generation by incorporating multiple aspects of quality assessment, thus providing a more nuanced and holistic evaluation framework.
@@ 5 Experiments
Table 2: Automatic evaluation results of existing interleaved generation models on InterleavedBench using InterleavedEval based on GPT-4o. TIC is the abbreviation for ’Text-Image Coherence’. The best results are highlighted in bold.
@@ 5.2 Main Results
We show the main results of using InterleavedEval to conduct the fine-grained evaluation for various baseline approaches on InterleavedBench in Table 2. The baselines in the upper part are the integrated and open-sourced models while the baselines in the lower part are the pipeline models where the LMMs are proprietary. From Table 2, we observe that: First, the pipeline models consistently outperform the integrated models on all evaluation aspects by a significant margin, where GPT-4o + DALL·E 3 achieves the best performance on helpfulness and the average score of all the aspects. This indicates that building a strong interleaved generation model for general purposes remains a significant challenge. Second, the pipeline models achieve significantly good performance on text quality since Gemini and GPT-4o have strong text generation capabilities. Also, the generated visual prompts are generally coherent with the text content and they are directly fed into the image generation model, so the performance on text-image coherence of pipeline models is also remarkable. Third, we observe that the common errors of integrated models include the output text and/or images being empty, in poor quality, or having severe duplication. This is probably due to their weak instruction-following abilities. Fourth, image coherence is the most challenging aspect for the pipeline models. This is because the image generation model cannot take the images in the input context or previously generated images as conditions. Thus, the generated images do not have strong coherence.
Given the closed-source nature of GPT-4o, the evaluation based on GPT-4o can be less transparent and sometimes may not be fully reproducible. To this end, we also implement our InterleavedEval using the current state-of-the-art open-sourced LMM, i.e., LLaVA-NeXT-Interleave Li et al. (2024b), which supports interleaved text and image inputs. We report the results in Table 10 in Appendix C.
@@ Results
Figure 6: Radar figures of evaluation results on each evaluation aspect for each task.
@@
To validate the effectiveness of our proposed metric, we conduct a correlation analysis by comparing the evaluation results from automatic metrics with our human evaluation results. Since the baseline metrics only predict an overall score for each instance, we use the same set of evaluation scores to compare against the human rating on each aspect separately. For InterleavedEval, we compare evaluation scores with the human rating on corresponding aspects. Since most baselines require a gold reference, we use the context-based subset, where each instance has an associated reference output, to compute the correlation. From Table 4, our InterleavedEval consistently outperforms previous metrics by a significant margin in every aspect. Our metric has a particularly higher correlation on text quality, which is because text quality is relatively easier to evaluate with large language models like GPT-4o Zheng et al. (2023b). Our metric achieves the lowest correlation on perceptual quality. The plausible reason is that GPT-4o’s perceptual recognition capability is still not strong enough to accurately detect visual artifacts or unnatural disruptions in the images Fu et al. (2024b). We also find that baseline metrics generally achieve poorer correlation, e.g., most metrics except for BERTScore almost do not have any correlation with helpfulness. BERTScore achieves the best correlation on helpfulness among baseline metrics, which indicates that text quality could be a good indicator of whether the overall interleaved content is helpful.
In addition, we also report the correlation with human judgments of InterleavedEval based on the open-sourced LLaVA-NeXT-Interleave in Table 4. InterleavedEval-LLaVA can achieve promising correlations with humans, generally surpassing previous metrics by a large margin. While there is still a significant gap between the performance of GPT-4o and LLaVA, probably due to the limited capability of LLaVA-NeXT-Interleave, we believe our InterleavedEval-LLaVA is a good alternative to InterleavedEval-GPT-4o in the scenarios where transparency and reproducibility are highly desired. We leave how to build a more powerful open-sourced evaluator for future work.
@@ 6 Discussions
MiniGPT5 1.29 3.47 2.04 2.64 1.76 2.24
GILL 1.37 3.96 2.01 2.61 1.51 2.29
EMU-2 1.29 2.22 1.65 1.18 1.84 1.64
Gemini1.5+SDXL 3.29 4.24 3.26 3.94 3.25 3.60
GPT-4o+DALLE3 3.12 4.39 3.08 4.36 3.48 3.69
Table 5: Automatic evaluation results of the context-based subset on InterleavedBench. TIC is the abbreviation for ’Text-Image Coherence’. The best results are highlighted in bold.
MiniGPT5 1.00 1.09 1.07 1.06 1.78 1.20
GILL 0.12 2.23 2.58 0.23 1.45 1.32
EMU-2 0.77 2.35 2.20 1.05 1.38 1.55
Gemini1.5+SDXL 4.50 3.66 4.13 3.98 4.10 4.07
GPT-4o+DALLE3 4.60 4.31 4.05 4.52 4.41 4.38
Table 6: Automatic evaluation results of the context-free subset on InterleavedBench. TIC is the abbreviation for ’Text-Image Coherence’. The best results are highlighted in bold.
Output Steps Text Quality Perceptual Quality Image Coherence TIC Helpfulness AVG
Less 1.8 1.1 1.2 1.3 2.1 1.5
Equal 2.7 3.8 4.0 4.0 3.0 3.5
More 1.7 3.5 2.4 3.3 2.0 2.6
Table 7: Analysis of the number of output steps compared with ground truths.
@@
We conduct a qualitative analysis of benchmarked models in Figure 5 and have the following observations: (1) while GILL can generate images with reasonable quality, the generated text and images are typically not coherent with the instruction and context. In the example in the first row, the generated text is totally irrelevant to the task, while the image is also inconsistent with input images. (2) EMU-2 can often generate text that is relevant to the task, but the quality is not good enough. In the example in the second row, it repeatedly says “soak the fabric in water” but does not contain other useful content. Another weakness of EMU-2 is its poor conditional image generation capability, where generated images have obvious visual distortions and could be duplicated with input images. (3) On the other hand, the pipeline models can generally better follow the instructions and generate text and images in higher quality. Nevertheless, they still occasionally have some drawbacks. For Gemini+SDXL, some of the generated images (e.g., the first output image in the second example) still have obvious defects. For GPT-4+DALLE3, the style of generated images can be dramatically different from input images, as DALLE3 is prone to generate images in cartoon or dramatic styles. (4) Maintaining image coherence, i.e., the coherence of style and entities across images, is still very challenging for most models. In the third example, for the pipeline models, the same character has a very different appearance across the images, which makes the content inconsistent. (5) For the instances on the context-free subset, the integrated baselines have significantly worse performance, where they only generate one image with extremely poor quality. We hypothesize the reason is that those models cannot truly understand and follow the instructions. To sum up, our qualitative analysis indicates there is still significant room for improvement in interleaved generation.
@@
Breakdown Performance on Context-based and Context-free Subsets
We show the breakdown performance on the context-based and context-free subsets of InterleavedBench in Table 5 and Table 6. Our findings are: (1) pipeline baselines consistently outperform integrated baselines on both subsets; (2) pipeline baselines have better performance on the context-free subset than the context-based subset, while integrated baselines have better performance on the context-based subset than the context-free subset. Based on the results and our observations, we find the following reasons that could contribute to the discrepancy in performance: (1) pipeline approaches first generate the text along with captions with target images, which can be considered as a planning stage to provide the basis on what images should be generated, making generated interleaved content more useful and reasonable; (2) using separate models (LLMs for text generation and T2I models for image generation) facilitates the generation of high-quality content in each modality; (3) Existing integrated models may struggle with the context-free subset because they haven’t been trained on data with text-only inputs and interleaved multimodal outputs.
Impact of the Number of Output Steps
We conduct an analysis of how the number of output steps affects the performance when compared with that in the ground truths. We calculate the performance of GPT4o-DALLE3 under three cases: the number of predicted steps is less, equal to, or larger than that in the ground truth (“Less”, “Equal”, “More”). From Table 7, when the number of predicted steps is less than the ground truths, the model performance is generally worse. This indicates that instances with fewer steps are considered as lower quality and less helpful. When the model has more output steps than ground truths, the performance on text quality, image coherence, and helpfulness are lower. This is because we observed the models produce more images than necessary. Often, these output images are repetitive of the input images or previously generated images. Since we explicitly penalize such repetition in our evaluation criteria, the performance for these cases is lower.
@@ 8 Limitation
Acknowledgement
This research is partially supported by a research award from Intuit AI Research, the award No. 2238940 from the Faculty Early Career Development Program (CAREER) of the National Science Foundation (NSF), and the U.S. DARPA ECOLE Program #HR001122S0052. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix B More Details on Evaluation
B.1 Human Evaluation
More Details on Human Evaluation Setup
Inter-Annotator Agreement
We show the IAA of our human evaluation in Table 9. The inter-annotator agreement is reasonably good. Note that the evaluation of interleaved generation is still quite subjective, open-ended, and challenging, even with our carefully designed human evaluation aspects and guidelines.
@@ B.1 Human Evaluation
Appendix C Additional Experiment Results
C.1 Automatic Evaluation Results on based on LLaVA-NeXT-Interleave
We report the automatic evaluation results in Table 10 based on the existing state-of-the-art open-sourced LMM that supports interleaved text and image inputs, i.e., LLaVA-NeXT-Interleaved Li et al. (2024b), in Table 10. We use the same evaluation instructions and criteria to prompt the model to predict numerical scores from 1 to 5. We show the automatic evaluation results in Table A and the correlation analysis in Table B. We use the same experiment setup for a fair comparison. From Table 10, the benchmarked performance using InterleavedEval with LlaVA-NeXT-Interleaved generally aligns with human evaluation in Table 3 and automatic evaluation with GPT-4o in Table 2. For example, the pipeline-based models consistently outperformed the integrated baselines, and GPT-4o-DALLE3 remains the best model overall.
MiniGPT-5 2.52 2.22 2.28 1.68 2.59 2.26
GILL 1.60 3.26 3.09 1.50 3.08 2.51
EMU-2 2.86 2.41 2.44 1.66 3.11 2.50
EMU-2 (Gold Text) 1.44 3.31 3.30 1.51 3.25 2.56
Gemini1.5+SDXL 3.70 3.86 3.79 3.73 3.78 3.77
GPT-4o+DALLE3 3.61 4.16 3.93 3.82 3.87 3.88
Table 10: Automatic evaluation results of existing interleaved generation models on InterleavedBench using InterleavedEval based on open-sourced LLaVA-NeXT-Interleave. TIC is the abbreviation for ’Text-Image Coherence’. The best results are highlighted in bold.